After reading Elizabeth Warren’s book, All Your Worth, you decide to be very careful about your...

  1. After reading Elizabeth Warren’s book, All Your Worth, you decide to be very careful about your LARGE purchases and commitments. As a result, you consistently choose a rent rate or a mortgage amount that is $100 less than you “could” afford. You also choose to pay cash for your cars to avoid $400 or more in interest expenses to finance a car purchase. In total, these choices save you $500 per month. How much can you accumulate if you save this $500 per month in an index fund that averages 9% per year for the 45 years you have until retirement? (Assume monthly compounding.)
Answer #1

Annual return on index fund =9%

Monthly return =(9/12)%=0.75%=0.0075

Number of months of savings =45*12=540

Compount Amount Factor (CAF)=(F/A,i,N)=(((1+i)^N)-1)/i

i=Monthly Return =0.75%=0.0075

N=Number of months=540

CAF=(F/A,0.75%,540)=(((1+0.0075)^540)-1)/0.0075=7404.878469

Monthly Savings=$500

Amount accumulated=500*CAF=500*7404.8786469=$3,702,439

Amount accumulated $3,702,439